OpenStudy (anonymous):

what is the period of the function 2csc (2πx)

OpenStudy (anonymous):

is the period 2pi/2?

OpenStudy (kainui):

A period is the time it takes to go through one complete cycle. You go through one cycle every 2pi radians. So if we look at the function, we can see that at csc(2pi) x=1. That means at 1, 2, 3, etc... you go through a complete cycle and your period is 1.

OpenStudy (kainui):

 See how we start back where we begin after 2pi? halfway around is pi. Kind of strange to get used to.

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

wait...i still dont get it isn't there a rule to figure out its 1? like

OpenStudy (kainui):

Wrong @Hr.Tboy http://www.wolframalpha.com/input/?i=period+of+2csc+%282%CF%80x%29

OpenStudy (anonymous):

for example to find the period of sec(2pix) you do the formula 2pi/b

OpenStudy (kainui):

I mean, I suppose there is probably a formula for it to figure it out just like there's a formula for area of a rectangle is A=l*w. But don't get lost in the formula, it's about the idea. A period occurs every time the function repeats itself. For instance, the period of the earth is 24 hours before it becomes the same time and repeats again.

OpenStudy (anonymous):

okay so in this question, 2csc (2πx), 2π is a period already so its 1

OpenStudy (anonymous):

ah 2pi/ 2pi = 1 makes sense!

hartnn (hartnn):

just another way to look at it: the period T of \(\csc (\omega t)\) is \(\large T= \frac{2\pi}{\omega}\) here, \(\omega =2\pi\)

OpenStudy (anonymous):

THANK YOU:)

hartnn (hartnn):

same goes for \(\sin(\omega t) \text{or any other trigonometric function..} \)

OpenStudy (kainui):

Thanks hartnn, sorry I couldn't write it out in an equation like that, it's a bit late here! =D

OpenStudy (anonymous):

@hartnn @kainui @isugano thnx all.. i learned too;)